One of the new deposit construction of our competitor credit bank offers 12% annual nominal interest rate but pays all interest semi-annually. Our bank wants to create a deposit, which pays quarterly, and its effective rate of return is 1% higher than the rival's effective rate of return. What nominal interest rate should we determine for the new product?
Answer #1

Credit Bank Annual Nominal Rate of interest =12%=0.12

Semi annual interest rate =6%=0.06

Effective annual interest rate =R

1+R=(1+0.06)^2=1.1236

Effective annual interest rate =1.1236-1=0.1236=12.36%

Effective rate of return of our bank =R+1%=12.36+1=13.36%=0.1336

Assume Quarterly interest rate of our bank =i

(1+i)^4=1+0.1336=1.1336

1+i=1.1336^(1/4)=1.031846

Quarterly interest rate =0.031846

Annual Nominal interest rate for the new product =0.031846*4=0.1274=12.74%

Annual Nominal interest rate for the new product 12.74%